OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y+3=-5(x-2) y-2=-5(x+3) y+3=-1/5(x-2) y-2=-1/5(x+3)

OpenStudy (anonymous):

Your first step would be to find the slope. Do you know how?

OpenStudy (anonymous):

y2-y1/ x2-x2?

OpenStudy (anonymous):

x1*

OpenStudy (anonymous):

Yep. Then you can plug the slope into m in the point-slope equation: \[y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

Wait... huh?

OpenStudy (anonymous):

so whatever you get as the slope goes into the "m" position. You need one point, either (-3,2) OR (2,1) for the y1 and x1 values

OpenStudy (anonymous):

I got 4 for the slope.

OpenStudy (anonymous):

OK, now use (-3,2) for the x1 and y1 values. It looks like the answer will be based on that point rather than (2,1) (fyi, but when you simplify them they're the same) .

OpenStudy (anonymous):

y-1=4(x-2)?

OpenStudy (anonymous):

yeah, that's right but looks like the answer choices want you to use (-3,2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y-2=m(x--3) which is y-2=m(x+3) might want to check your slope.

OpenStudy (anonymous):

...... I'm so lost.

OpenStudy (anonymous):

sorry! :( [slope]=m=(1-2)/(2--3)=(-1)/(2+3)=-1/5

OpenStudy (anonymous):

So slope= -5

OpenStudy (anonymous):

no slope is -1/5

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

sorry for misleading you by not checking it myself earlier ~ so now, using (-3,2) as (x1,y1)\[y-2=-{1 \over 5}(x+3)\]

OpenStudy (anonymous):

It's fine. Thank you for he help

OpenStudy (anonymous):

sorry it took so long. hope it helped at least a little.